The aqueous solution contains 34.0% H3PO4 by weight and has a density of 1,209 g / ml. What is the molarity of this solution?


1
Expert's answer
2021-01-15T06:26:49-0500

So in 100 g solution 34 g H3PO4 present

Molar mass of H3PO4 = (3×1) +(31×1) +(16×4) = 3+31+64 = 98 gmol-1

Mol of H3PO4 = 34\98 = 0.347 mol

100 g solution has a volume = 100\1.209 = 82.71 ml as volume = mass\ density

So molarity of solution

= (mol of solute)×1000 \(volume of solution in ml)

= (0.347×1000) \82.71

=4.19 M (ans)